Question provided by co-worker I'm trying to help...
Given the equation
x^2 + 6sqrt2x + y^2 - 4sqrt5y = 3
complete the square for both x and y and find the equation of the circle.
A. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 13
B. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 35
C. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 38
D. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 10
I can usually help her figure things out, or we at least muddle through them together. It's been so long, though, that I'm lost on this on.
I thank you, and she thanks you. Answer by stanbon(75887) 
----------------------------------
complete the square for both x and y and find the equation of the circle.
A. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 13
B. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 35
C. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 38
D. (x + 3sqrt2)^2 + (y - 2sqrt5)^2 = 10
-----------------------------------------------
Comment: The difference between the answer to your posted problem
and the answer "c" (above) is the 3 in the posted problem.
Are you sure that number is not -3 ?
===============================
Cheers,
Stan H.